# Tiered Reward Functions: Specifying and Fast Learning of Desired Behavior

### v2

@@ Appendix D Tabular Grid Worlds Experiment Details
The implementation of many tabular environments and algorithms is based on the MSDM library by Ho, Correa, and Ritter (2021). All MSDM experiments are run on an Ubuntu $18.04$ system with an Intel Core i7-9700K CPU and $32$G of RAM.
Following Koenig and Simmons (1993), we use a greedy policy for action selection and initialize the Q-values optimistically to exploit directed exploration. Since all reward values are non-positive, it is sufficient to initialize the Q-values to 00. We use a learning rate of $\alpha=0.90$ (tuned from $\alpha\in\{0.95,0.90,0.85\}$, all of which performed similarly). We set the small constant $\delta=0.1$.
Appendix E MiniGrid Experiment Details
@@
The PPO agent used an ImpalaCNN architecture for the policy and value network. The architecture of the policy network is as follows:
@@ 1 $-1$ $-1$
Appendix I Additional Figures

### v3

@@ Appendix D Tabular Grid Worlds Experiment Details
The implementation of many tabular environments and algorithms is based on the MSDM library by Ho et al. (2021). All MSDM experiments are run on an Ubuntu $18.04$ system with an Intel Core i7-9700K CPU and $32$G of RAM.
Following Koenig & Simmons (1993), we use a greedy policy for action selection and initialize the Q-values optimistically to exploit directed exploration. Since all reward values are non-positive, it is sufficient to initialize the Q-values to 00. We use a learning rate of $\alpha=0.90$ (tuned from $\alpha\in\{0.95,0.90,0.85\}$, all of which performed similarly). We set the small constant $\delta=0.1$.
Appendix E RMAX results
Appendix F MiniGrid Experiment Details
@@
The PPO agent used an ImpalaCNN architecture (Espeholt et al., 2018) for the policy and value network. The architecture of the policy network is as follows:
@@ 1 $-1$ $-1$
Appendix I Additional Figure: Pareto-optimal Rewards